# Stability-Certified On-Policy Data-Driven LQR via Recursive Learning and Policy Gradient

### v1

@@ Appendix A Preliminaries on averaging theory for two-time-scale systems
We note that (29) is obtained by setting $K_{t}=K^{\star}$ in (4.1) (which compactly collects the updates (19a), (19b), and (4.1)). Hence, we start by inspecting (19a) and (19b) restricted to the manifold in which $K_{t}=K^{\star}$, namely
@@
System (B) is a cascade, therefore its steady-state solution is $\operatorname{col}(\mathrm{w}_{t},\mathrm{\mathrm{x}}_{t})=\operatorname{col}(I_{n_{\mathrm{w}}},\Pi_{\mathrm{x}})\mathrm{w}_{t}$, with $\Pi_{\mathrm{x}}\in\mathbb{R}^{n\times m}$ solution to the following Sylvester equation
$\Pi_{\mathrm{x}}F=({A}_{\star}+{B}_{\star}K^{\star})\Pi_{\mathrm{x}}+{B}_{\star}E.$ (B.2)
Being $F$ marginally stable (cf. Assumption 3.1) and ${A}_{\star}+{B}_{\star}K^{\star}$ Schur (so that $\sigma(F)\cap\sigma({A}_{\star}+{B}_{\star}K^{\star})=\emptyset$) the solution $\Pi_{\mathrm{x}}$ exists and is unique. Then, we inspect the dynamics (4.1) restricted to the manifold in which $\mathrm{\mathrm{x}}_{t}=\Pi_{\mathrm{x}}\mathrm{w}_{t}$ and $K_{t}=K^{\star}$. Let $M\in\mathbb{R}^{(n+m)\times(n+m)}$ be
$M:=\begin{bmatrix}\Pi_{\mathrm{x}}^{\top}&(K^{\star}\Pi_{\mathrm{x}}+E)^{\top}\end{bmatrix}^{\top},$ (B.4)
then it holds
$\mathtt{vec}\left(\mathrm{w}_{{t+1}}\mathrm{w}_{{t+1}}^{\top}\right)=\mathtt{vec}\left(F\mathrm{w}_{t}\mathrm{w}_{t}^{\top}F^{\top}\right)$ (B.5a)
$H^{\text{vc}}_{t+1}=\lambda H^{\text{vc}}_{t}+\mathtt{vec}\left(M\mathrm{w}_{t}\mathrm{w}_{t}^{\top}M^{\top}\right)$ (B.5b)
$S^{\text{vc}}_{t+1}=\lambda S^{\text{vc}}_{t}+\mathtt{vec}\left(M\mathrm{w}_{t}\mathrm{w}_{t}^{\top}M^{\top}\theta^{\star}\right),$ (B.5c)
where the first equation comes from the vectorization of (19a). By exploiting the vectorization properties [^2] Given any $X_{1}\in\mathbb{R}^{n_{1}\times n_{2}}$, $X_{2}\in\mathbb{R}^{n_{2}\times n_{3}}$, and $X_{3}\in\mathbb{R}^{n_{3}\times n_{4}}$, it holds $\mathtt{vec}\left(X_{1}X_{2}X_{3}\right)=(X_{3}^{\top}\otimes X_{1})\mathtt{vec}\left(X_{2}\right)$., we can manipulate (B) to obtain the system
$\mathtt{vec}\!\left(\mathrm{w}_{{t+1}}\mathrm{w}_{{t+1}}^{\top}\right)\!=\!(F\otimes F)\mathtt{vec}\left(\mathrm{w}_{t}\mathrm{w}_{t}^{\top}\right)$ (B.6a)
$H^{\text{vc}}_{t+1}\!=\!\lambda H^{\text{vc}}_{t}\!+\!(M\otimes M)\mathtt{vec}\left(\mathrm{w}_{t}\mathrm{w}_{t}^{\top}\right)$ (B.6b)
$S^{\text{vc}}_{t+1}\!=\!\lambda S^{\text{vc}}_{t}\!+\!({\theta^{\star}}^{\top}\!M\!\otimes\!M)\mathtt{vec}\left(\mathrm{w}_{t}\mathrm{w}_{t}^{\top}\right)\!,$ (B.6c)
which enjoys again a cascade structure. Thus, let $\Pi_{H}\in\mathbb{R}^{(n+m)^{2}\times n_{\mathrm{w}}^{2}}$ and $\Pi_{S}\in\mathbb{R}^{(n+m)n\times n_{\mathrm{w}}^{2}}$ be the solution to the Sylvester equations associated to the cascade in (B) given by
$\Pi_{H}(F\otimes F)$ $=\lambda\Pi_{H}+M\otimes M$ (B.7a)
$\Pi_{S}(F\otimes F)$ $=\lambda\Pi_{S}+({\theta^{\star}}^{\top}M)\otimes M.$ (B.7b)
Being $F$ marginally stable (cf. Assumption 3.1) and $\lambda\in(0,1)$, then $\sigma(F\otimes F)\cap\sigma(\lambda I)=\emptyset$ and, thus, the solutions $\Pi_{S}$ and $\Pi_{H}$ to (B) exist and are unique. The proof of (29) follows by (i) noticing that $(\Pi_{\mathrm{x}},\Pi_{H},\Pi_{S})$ are used to define $\chi^{{\textnormal{ss}}}$ (cf. (28)), and (ii) plugging (B.2) and (B) into system (4.1).
As for (30), it can be shown using an algebraic manipulation of (B). Indeed, by pre-multiplying (B.7a) by ${\theta^{\star}}^{\top}\otimes I_{n+m}$ we can write
$({\theta^{\star}}^{\top}\!\otimes\!I_{n+m})\Pi_{H}(F\otimes F-\lambda I)$ $\!=\!({\theta^{\star}}^{\top}\otimes I_{n+m})(M\otimes M)$
=$\!\stackrel{{\scriptstyle(a)}}{{=}}\!({\theta^{\star}}^{\top}M)\otimes M$
=$\!\stackrel{{\scriptstyle(b)}}{{=}}\!\!\Pi_{S}(\!F\!\otimes\!F\!-\!\lambda I),$ (B.8)
where in $(a)$ we used the mixed-product property of the Kronecker operator [^3] Given any $X_{1}\!\in\!\mathbb{R}^{n_{1}\times n_{2}}$, $X_{2}\!\in\!\mathbb{R}^{n_{3}\times n_{4}}$, $X_{3}\!\in\!\mathbb{R}^{n_{2}\times n_{5}}$, and $X_{4}\!\in\!\mathbb{R}^{n_{4}\times n_{6}}$, it holds $(X_{1}\!\otimes\!X_{2})(X_{3}\!\otimes\!X_{4})\!=\!(X_{1}X_{3})\!\otimes\!(X_{2}X_{4})$., while $(b)$ follows from (B.7b). So that the proof is complete.
Appendix C Proof of Lemma 4.2
Lemma C.1
Let the assumptions of Theorem 3.2 hold true. Then, the matrix $H^{{\textnormal{ss}}}_{t}$, defined in (34), is invertible for all $t\in\mathbb{N}$. normal-$\blacksquare$
Proof C.2
We will prove the invertibility property of the matrix $H^{{\textnormal{ss}}}_{t}$ by investigating the evolution of $H_{t}$. The dynamics of $H_{t}$ in (19g) restricted to the manifold in which $\mathrm{\mathrm{x}}_{t}=\Pi_{\mathrm{x}}\mathrm{w}_{t}$ and $K_{t}=K^{\star}$ reads as
$H_{{t+1}}=\lambda H_{t}+M\mathrm{w}_{t}\mathrm{w}_{t}^{\top}M^{\top},$ (C.1)
with $M$ as in (B.4). The explicit solution of (C.1) is
$H_{t}=\lambda^{t}H_{0}+M\underbrace{\left(\sum_{\tau=0}^{t-1}\lambda^{t-1-\tau}\mathrm{w}_{\tau}\mathrm{w}_{\tau}^{\top}\right)}_{\mathcal{W}_{t}}M^{\top}.$ (C.2)
Being $\lambda\in(0,1)$, the free evolution $\lambda^{t}H_{0}$ in (C.2) vanishes as $t\to\infty$. Hence, it does not impact on the invertibility of the steady-state solution.
Therefore, let us focus on the forced response $M\mathcal{W}_{t}M^{\top}$ only. We first notice that $\alpha_{1}I_{n_{\mathrm{w}}}\leq\sum_{\tau=\bar{t}+1}^{\bar{t}+t_{\mathrm{w}}}\mathrm{w}_{\tau}\mathrm{w}_{\tau}^{\top}\leq\alpha_{2}I_{n_{\mathrm{w}}}$ for all $\bar{t}\in\mathbb{N}$ (cf. Assumption 3.1). Hence we can invoke [55, Lemma 1] to assert the positive definiteness of $\mathcal{W}_{t}$ for all $t\geq t_{\mathrm{w}}$. Let us consider the Cholesky decomposition of $\mathcal{W}_{t}$ given by $\mathcal{W}_{t}=\mathcal{C}_{t}\mathcal{C}_{t}^{\top}$, with $\mathcal{C}_{t}\in\mathbb{R}^{n_{\mathrm{w}}\times n_{\mathrm{w}}}$ invertible. Then [^4] Given any $X\!\in\!\mathbb{R}^{n\times m}$, it holds $\operatorname{rank}(XX^{\top})\!=\!\operatorname{rank}(X)\!=\!\operatorname{rank}(X^{\top})$., for all $t\geq t_{\mathrm{w}}$, we can write
$\operatorname{rank}\left(M\mathcal{W}_{t}M^{\top}\right)$ $=\operatorname{rank}(M\mathcal{C}_{t})\stackrel{{\scriptstyle(a)}}{{=}}\operatorname{rank}(M),$ (C.3)
$\mathrm{w}_{{t+1}}$ $=F\mathrm{w}_{t}$ (C.4a)
$\mathrm{\mathrm{x}}_{{t+1}}$ $=({A}_{\star}+{B}_{\star}K^{\star})\mathrm{\mathrm{x}}_{t}+{B}_{\star}E\mathrm{w}_{t}.$ (C.4b)
@@
Recalling that $\mathrm{d}_{t}=E\mathrm{w}_{t}$ satisfies condition (18f) (cf. Assumption 3.1) and that $({A}_{\star},{B}_{\star})$ is controllable (cf. Assumption 2.1), we can invoke [48, Cor. 2] to claim that
@@
$\operatorname{rank}\left(\begin{bmatrix}\mathrm{x}_{0}&\dots&\mathrm{x}_{\mathbf{t}_{\mathrm{d}}-1}\\ \mathrm{d}_{0}&\dots&\mathrm{d}_{\mathbf{t}_{\mathrm{d}}-1}\end{bmatrix}\right)=n+m,$ (C.7)
@@
for all $(\mathrm{x}_{0},\mathrm{d}_{0})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$. When the initial condition of (C.4b) lies in the invariant steady-state locus (cf. (B.2)), i.e., when $\mathrm{x}_{0}=\Pi_{\mathrm{x}}\mathrm{w}_{0}$, the condition in (C.7) simplifies to
@@
$\operatorname{rank}\left(M\begin{bmatrix}\mathrm{w}_{0}&\dots&\mathrm{w}_{\mathbf{t}_{\mathrm{d}}-1}\end{bmatrix}\right)=n+m,$ (C.9)
@@
with $M$ as in (B.4), which allows us to conclude that [^6] Given $X_{1}\in\mathbb{R}^{n_{1}\times n_{2}}$ and $X_{2}\in\mathbb{R}^{n_{2}\times n_{3}}$, it holds $\operatorname{rank}(X_{1}X_{2})\leq\min\{\operatorname{rank}(X_{1}),\operatorname{rank}(X_{2})\}$.
@@
Moreover, being $\operatorname{rank}(M)\leq n+m$ by construction, the above inequality yields to $\operatorname{rank}(M)=n+m$, which, in turn, combined with (C.3), allows us to write
$\operatorname{rank}\left(M\mathcal{W}_{t}M^{\top}\right)=n+m,$ (C.10)
@@
for all $t\geq t_{\infty}$. Finally, being $H^{{\textnormal{ss}}}_{t}$ a static function of the periodic signal $\mathrm{w}_{t}$, then $H^{{\textnormal{ss}}}_{t}$ is periodic as well so that its full-rankness is independent of $t$. Thus, it must be that $\operatorname{rank}(H^{{\textnormal{ss}}}_{t})=n+m$ for all $t\in\mathbb{N}$, and the proof follows. normal-$\blacksquare$
Once the invertibility of $H^{{\textnormal{ss}}}_{t}$ has been established by Lemma C.1, we are ready to prove Lemma 4.2. Let us label the two components of $f^{\textsc{av}}$ as
$\begin{bmatrix}f^{\textsc{av}}_{1}(\tilde{\mathrm{z}})\\ f^{\textsc{av}}_{2}(\tilde{\mathrm{z}})\end{bmatrix}:=\lim_{T\to\infty}\frac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}f(0,\tilde{\mathrm{z}},\tau).$ (C.13)
As for $f^{\textsc{av}}_{1}(\mathrm{z})$, we can write
$f^{\textsc{av}}_{1}(\tilde{\mathrm{z}})$ $=-\lim_{T\to\infty}\frac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}(H^{{\textnormal{ss}}}_{\tau})^{\dagger}H^{{\textnormal{ss}}}_{\tau}\tilde{\theta}\stackrel{{\scriptstyle(a)}}{{=}}-\tilde{\theta},$
where in $(a)$ we used Lemma C.1 to guarantee the invertibility of $H^{{\textnormal{ss}}}_{t}$ for all $t\in\mathbb{N}$ As for $f^{\textsc{av}}_{2}(\mathrm{z})$, its existence is trivially shown by observing that it does not depend on $t$. Hence, given $\tilde{\mathrm{z}}=\operatorname{col}(\tilde{K},\tilde{\theta})$, it holds
$f^{\textsc{av}}_{2}(\tilde{\mathrm{z}}):=-G(\tilde{K}+K^{\star},\tilde{\theta}+\theta^{\star}).$
Appendix D Proof of Proposition 4.3
@@
To ease the notation, we start by decomposing the state of (40) as $\tilde{\mathrm{z}}^{\textsc{av}}_{t}:=\operatorname{col}(\tilde{\theta}^{\textsc{av}}_{t},\tilde{K}^{\textsc{av}}_{t})$. Then, we recall [30, Lemma 3.12] to guarantee that the cost $J$, defined in (6), is gradient dominated, that is for all $K\in\mathcal{D}$ it holds
$J(K,\theta^{\star})-J(K^{\star},\theta^{\star})\leq\mu\left\|G(K,\theta^{\star})\right\|^{2},$ (D.1)
@@
$V(\tilde{K}^{\textsc{av}},\tilde{\theta}^{\textsc{av}}):=$ (D.2)
$\hskip 22.76228pt\kappa\left(J(\tilde{K}^{\textsc{av}}+K^{\star},\theta^{\star})-J(K^{\star},\theta^{\star})\right)+\tfrac{1}{2}\left\|\tilde{\theta}^{\textsc{av}}\right\|^{2},$
@@
Let $c_{0}>0$ be the smallest number such that $(\tilde{K}^{\textsc{av}}_{0},\tilde{\theta}^{\textsc{av}}_{0})\in\Omega_{c_{0}}$ and define
$\beta_{1}\!$ $:=\!\!\max_{(\tilde{K}^{\textsc{av}},\tilde{\theta}^{\textsc{av}})\in\Omega_{c_{0}}}\!\!\left\{\left\|\dfrac{\partial G(\tilde{K}^{\textsc{av}}\!+\!K^{\star},\tilde{\theta}^{\textsc{av}}\!+\!\theta^{\star})}{\partial\tilde{K}^{\textsc{av}}}\right\|\right\}$ (D.3a)
$\beta_{2}\!$ $:=\!\!\max_{(\tilde{K}^{\textsc{av}},\tilde{\theta}^{\textsc{av}})\in\Omega_{c_{0}}}\!\!\left\{\left\|\dfrac{\partial G(\tilde{K}^{\textsc{av}}\!+\!K^{\star},\tilde{\theta}^{\textsc{av}}\!+\!\theta^{\star})}{\partial\tilde{\theta}^{\textsc{av}}}\right\|\right\}\!\!.$ (D.3b)
@@
$=\kappa\Big{(}J(\tilde{K}^{\textsc{av}}_{{t+1}}+K^{\star},\theta^{\star})-J(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\Big{)}$
$\hskip 11.38092pt-\gamma\left(1-\gamma/2\right)\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}$
@@
$\hskip 11.38092pt-\kappa J(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})-\gamma\left(1-\gamma/2\right)\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\kappa J(\tilde{K}^{\textsc{av}}_{t}+K^{\star}-\gamma G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\tilde{\theta}^{\textsc{av}}_{t}+\theta^{\star}),\theta^{\star})$
$\hskip 11.38092pt-\gamma\kappa\left(1-\gamma\tfrac{\beta_{1}}{2}\right)\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2}$
$\hskip 11.38092pt-\gamma\left(1-\gamma/2\right)\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}\!,$ (D.4)
where $(a)$ uses the update of $\tilde{K}^{\textsc{av}}_{{t+1}}$ and adds ±$\pm\kappa J(\tilde{K}^{\textsc{av}}_{t}+K^{\star}-\gamma G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star}),\theta^{\star})$, $(b)$ uses the Taylor expansion of $J(\cdot,\cdot)$ about $(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})$ evaluated at $(\tilde{K}^{\textsc{av}}_{t}+K^{\star}-\gamma G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star}),\theta^{\star})$ and uses (D.3a).
Next, we manipulate the difference between the first two terms in (D.4). By expanding $J(\cdot,\cdot)$ about $(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})$ evaluated at $(\tilde{K}^{\textsc{av}}_{t}+K^{\star}-\gamma G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\tilde{\theta}^{\textsc{av}}_{t}+\theta^{\star}),\theta^{\star})$ and $(\tilde{K}^{\textsc{av}}_{t}+K^{\star}-\gamma G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star}),\theta^{\star})$ and using (D.3a) and the Cauchy-Schwarz inequality, we can write

### v2

@@ 6 Conclusions
Lemma B.1
Let the assumptions of Theorem 3.3 hold true. Then, the matrix $H^{{\textnormal{ss}}}_{t}$, defined in (32), is invertible for all $t\in\mathbb{N}$.
Proof B.2
We will prove the invertibility property of the matrix $H^{{\textnormal{ss}}}_{t}$ by investigating the evolution of $H_{t}$. The dynamics of $H_{t}$ in (19c) restricted to the manifold in which $\mathrm{\mathrm{x}}_{t}=\Pi_{\mathrm{x}}\mathrm{w}_{t}$ and $K_{t}=K^{\star}$ reads as
$H_{{t+1}}=\lambda H_{t}+M\mathrm{w}_{t}\mathrm{w}_{t}^{\top}M^{\top},$ (B.1)
with $M$ as in (A.3). The explicit solution of (B.1) is
$H_{t}=\lambda^{t}H_{0}+M\underbrace{\left(\sum_{\tau=0}^{t-1}\lambda^{t-1-\tau}\mathrm{w}_{\tau}\mathrm{w}_{\tau}^{\top}\right)}_{\mathcal{W}_{t}}M^{\top}.$ (B.2)
Being $\lambda\in(0,1)$, the free evolution $\lambda^{t}H_{0}$ in (B.2) vanishes as $t\to\infty$. Hence, it does not impact on the invertibility of the steady-state solution.
Therefore, let us focus on the forced response $M\mathcal{W}_{t}M^{\top}$ only. We first notice that $\alpha_{1}I_{n_{\mathrm{w}}}\leq\sum_{\tau=\bar{t}+1}^{\bar{t}+t_{\mathrm{w}}}\mathrm{w}_{\tau}\mathrm{w}_{\tau}^{\top}\leq\alpha_{2}I_{n_{\mathrm{w}}}$ for all $\bar{t}\in\mathbb{N}$ (cf. Assumption 3.1). Hence we can invoke [48, Lemma 1] to assert the positive definiteness of $\mathcal{W}_{t}$ for all $t\geq t_{\mathrm{w}}$. Let us consider the Cholesky decomposition of $\mathcal{W}_{t}$ given by $\mathcal{W}_{t}=\mathcal{C}_{t}\mathcal{C}_{t}^{\top}$, with $\mathcal{C}_{t}\in\mathbb{R}^{n_{\mathrm{w}}\times n_{\mathrm{w}}}$ invertible. Then [^4] Given any $X\!\in\!\mathbb{R}^{n\times m}$, it holds $\operatorname{rank}(XX^{\top})\!=\!\operatorname{rank}(X)\!=\!\operatorname{rank}(X^{\top})$., for all $t\geq t_{\mathrm{w}}$, we can write
$\operatorname{rank}\left(M\mathcal{W}_{t}M^{\top}\right)$ $=\operatorname{rank}(M\mathcal{C}_{t})\stackrel{{\scriptstyle(a)}}{{=}}\operatorname{rank}(M),$ (B.3)
@@
Recalling that $\mathrm{d}_{t}=E\mathrm{w}_{t}$ satisfies condition (18b) (cf. Assumption 3.1) and that $({A}_{\star},{B}_{\star})$ is controllable (cf. Assumption 2.7), we can invoke [50, Cor. 2] to claim that
@@
$\operatorname{rank}\left(\begin{bmatrix}\mathrm{x}_{0}&\dots&\mathrm{x}_{\mathbf{t}_{\mathrm{d}}-1}\\ \mathrm{d}_{0}&\dots&\mathrm{d}_{\mathbf{t}_{\mathrm{d}}-1}\end{bmatrix}\right)=n+m,$ (B.5)
@@
for all $(\mathrm{x}_{0},\mathrm{d}_{0})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$. When the initial condition of (B.4b) lies in the invariant steady-state locus (cf. (A.2)), i.e., when $\mathrm{x}_{0}=\Pi_{\mathrm{x}}\mathrm{w}_{0}$, the condition in (B.5) simplifies to
@@
$\operatorname{rank}\left(M\begin{bmatrix}\mathrm{w}_{0}&\dots&\mathrm{w}_{\mathbf{t}_{\mathrm{d}}-1}\end{bmatrix}\right)=n+m,$ (B.6)
@@
with $M$ as in (A.3), which allows us to conclude that [^6] Given $X_{1}\in\mathbb{R}^{n_{1}\times n_{2}}$ and $X_{2}\in\mathbb{R}^{n_{2}\times n_{3}}$, it holds $\operatorname{rank}(X_{1}X_{2})\leq\min\{\operatorname{rank}(X_{1}),\operatorname{rank}(X_{2})\}$.
@@
Moreover, being $\operatorname{rank}(M)\leq n+m$ by construction, the above inequality yields to $\operatorname{rank}(M)=n+m$, which, in turn, combined with (B.3), allows us to write
$\operatorname{rank}\left(M\mathcal{W}_{t}M^{\top}\right)=n+m,$ (B.7)
@@
for all $t\geq t_{\infty}$. Finally, being $H^{{\textnormal{ss}}}_{t}$ a static function of the periodic signal $\mathrm{w}_{t}$, then $H^{{\textnormal{ss}}}_{t}$ is periodic as well so that its full-rankness is independent of $t$. Thus, it must be that $\operatorname{rank}(H^{{\textnormal{ss}}}_{t})=n+m$ for all $t\in\mathbb{N}$, and the proof follows. $\blacksquare$
Once the invertibility of $H^{{\textnormal{ss}}}_{t}$ has been established by Lemma B.1, we are ready to prove Lemma 4.2. Let us label the two components of $f^{\textsc{av}}$ as
$\begin{bmatrix}f^{\textsc{av}}_{1}(\tilde{\mathrm{z}})\\ f^{\textsc{av}}_{2}(\tilde{\mathrm{z}})\end{bmatrix}:=\lim_{T\to\infty}\frac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}f(0,\tilde{\mathrm{z}},\tau).$
As for $f^{\textsc{av}}_{1}(\mathrm{z})$, its existence follows by the time independency of observing the first block of $f$ (cf. (31d)). Hence, it trivially coincides with $f_{1}$, namely, by denoting $\tilde{\mathrm{z}}=\operatorname{col}(\tilde{K},\tilde{\theta})$, we have
$f^{\textsc{av}}_{1}(\tilde{\mathrm{z}}):=-G(\tilde{K}+K^{\star},\tilde{\theta}+\theta^{\star}).$
As for $f^{\textsc{av}}_{2}(\mathrm{z})$, by using the expression of the second block of $f$ (cf. (31e)) restricted to $\tilde{\chi}=0$, we get
$f^{\textsc{av}}_{2}(\tilde{\mathrm{z}})$ $=-\lim_{T\to\infty}\frac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}(H^{{\textnormal{ss}}}_{\tau})^{\dagger}H^{{\textnormal{ss}}}_{\tau}\tilde{\theta}\stackrel{{\scriptstyle(a)}}{{=}}-\tilde{\theta},$
where in $(a)$ we exploited Lemma B.1 that guarantees the invertibility of $H^{{\textnormal{ss}}}_{t}$ for all $t\in\mathbb{N}$.
Appendix C Proof of Proposition 4.3
@@
To ease the notation, we start by decomposing the state of (36) as $\tilde{\mathrm{z}}^{\textsc{av}}_{t}:=\operatorname{col}(\tilde{K}^{\textsc{av}}_{t},\tilde{\theta}^{\textsc{av}}_{t})$. Then, we recall [30, Lemma 3.12] to guarantee that the cost $J$, defined in (9), is gradient dominated, that is for all $K\in\mathcal{D}$ it holds
$J(K,\theta^{\star})-J(K^{\star},\theta^{\star})\leq\mu\left\|G(K,\theta^{\star})\right\|^{2},$ (C.1)
@@
$V(\tilde{K}^{\textsc{av}},\tilde{\theta}^{\textsc{av}}):=\kappa\left(J(\tilde{K}^{\textsc{av}}+K^{\star},\theta^{\star})-J(K^{\star},\theta^{\star})\right)+\tfrac{1}{2}\left\|\tilde{\theta}^{\textsc{av}}\right\|^{2},$ (C.2)
@@
Let $c_{0}>0$ be the smallest number such that $(\tilde{K}^{\textsc{av}}_{0},\tilde{\theta}^{\textsc{av}}_{0})\in\Omega_{c_{0}}$ and introduce $\beta_{0},\beta_{1},\beta_{2}>0$ defined as
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta_{0}}$ $:={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\!\!\max_{(\tilde{K}^{\textsc{av}},\tilde{\theta}^{\textsc{av}})\in\Omega_{c_{0}}}\!\!\left\{\left\|G(\tilde{K}^{\textsc{av}}\!+\!K^{\star},\tilde{\theta}^{\textsc{av}}\!+\!\theta^{\star})\right\|\right\}}$ (C.3a)
$\beta_{1}\!$ $:=\!\!\max_{(\tilde{K}^{\textsc{av}},\tilde{\theta}^{\textsc{av}})\in\Omega_{c_{0}}}\!\!\left\{\left\|\dfrac{\partial G(\tilde{K}^{\textsc{av}}\!+\!K^{\star},\tilde{\theta}^{\textsc{av}}\!+\!\theta^{\star})}{\partial\tilde{K}^{\textsc{av}}}\right\|\right\}$ (C.3b)
$\beta_{2}\!$ $:=\!\!\max_{(\tilde{K}^{\textsc{av}},\tilde{\theta}^{\textsc{av}})\in\Omega_{c_{0}}}\!\!\left\{\left\|\dfrac{\partial G(\tilde{K}^{\textsc{av}}\!+\!K^{\star},\tilde{\theta}^{\textsc{av}}\!+\!\theta^{\star})}{\partial\tilde{\theta}^{\textsc{av}}}\right\|\right\}\!.$ (C.3c)
@@
$=\kappa\Big{(}J(\tilde{K}^{\textsc{av}}_{{t+1}}+K^{\star},\theta^{\star})-J(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\Big{)}-\gamma\left(1-\gamma/2\right)\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}$
@@
$\hskip 11.38092pt-\kappa J(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})-\gamma\left(1-\gamma/2\right)\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2},$ (C.4)
where $(a)$ uses the update of $\tilde{K}^{\textsc{av}}_{{t+1}}$ and adds ±$\pm\kappa J(\tilde{K}^{\textsc{av}}_{t}+K^{\star}-\gamma G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star}),\theta^{\star})$. Now, we recall that the set of stabilizing gains $\mathcal{K}$ is open [47, Lemma IV.3]. Then, since $\left\|G(\tilde{K}^{\textsc{av}}+K^{\star},\tilde{\theta}^{\textsc{av}}+\theta^{\star})\right\|$ is finite in $\Omega_{c_{0}}$ (cf. (C.3a)), we can arbitrarily fix a finite $\tilde{c}_{0}>c_{0}$ and find $\bar{\gamma}_{0}$ such that, for all $\gamma\in(0,\bar{\gamma}_{0})$, it holds
$(\tilde{K}^{\textsc{av}}+\gamma(G(\tilde{K}^{\textsc{av}}+K^{\star},\tilde{\theta}^{\textsc{av}}+\theta^{\star})),\tilde{\theta}^{\textsc{av}})\in\Omega_{c_{0}},$
for all $(\tilde{K}^{\textsc{av}},\tilde{\theta}^{\textsc{av}})\in\Omega_{c_{0}}$. To take advantage of this property, we introduce $\beta_{3},\beta_{4}>0$ defined as
$\beta_{3}\!$ $:=\!\!\max_{(\tilde{K}^{\textsc{av}},\tilde{\theta}^{\textsc{av}})\in\Omega_{\tilde{c}_{0}}}\!\!\left\{\left\|\dfrac{\partial G(\tilde{K}^{\textsc{av}}\!+\!K^{\star},\tilde{\theta}^{\textsc{av}}\!+\!\theta^{\star})}{\partial\tilde{K}^{\textsc{av}}}\right\|\right\}$ (C.5a)
$\beta_{4}\!$ $:=\!\!\max_{(\tilde{K}^{\textsc{av}},\tilde{\theta}^{\textsc{av}})\in\Omega_{\tilde{c}_{0}}}\!\!\left\{\left\|\dfrac{\partial G(\tilde{K}^{\textsc{av}}\!+\!K^{\star},\tilde{\theta}^{\textsc{av}}\!+\!\theta^{\star})}{\partial\tilde{\theta}^{\textsc{av}}}\right\|\right\}\!.$ (C.5b)
By using the Taylor expansion of $J(\cdot,\cdot)$ about $(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})$ evaluated at $(\tilde{K}^{\textsc{av}}_{t}+K^{\star}-\gamma G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star}),\theta^{\star})$ and using (C.5a) we can further bound (C.4) as
$\Delta V(\tilde{K}^{\textsc{av}}_{t},\tilde{\theta}^{\textsc{av}}_{t})$ $\leq\kappa J(\tilde{K}^{\textsc{av}}_{t}+K^{\star}-\gamma G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\tilde{\theta}^{\textsc{av}}_{t}+\theta^{\star}),\theta^{\star})$
$\hskip 11.38092pt-\kappa J(\tilde{K}^{\textsc{av}}_{t}+K^{\star}-\gamma G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star}),\theta^{\star})$ (C.6)
$\hskip 11.38092pt-\gamma\kappa\left(1-\gamma\tfrac{\beta_{3}}{2}\right)\left\|G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})\right\|^{2}-\gamma\left(1-\gamma/2\right)\left\|\tilde{\theta}^{\textsc{av}}_{t}\right\|^{2}.$
Next, we manipulate the difference between the first two terms in (C.6). By expanding $J(\cdot,\cdot)$ about $(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star})$ evaluated at $(\tilde{K}^{\textsc{av}}_{t}+K^{\star}-\gamma G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\tilde{\theta}^{\textsc{av}}_{t}+\theta^{\star}),\theta^{\star})$ and $(\tilde{K}^{\textsc{av}}_{t}+K^{\star}-\gamma G(\tilde{K}^{\textsc{av}}_{t}+K^{\star},\theta^{\star}),\theta^{\star})$ and using (C.5a) and the Cauchy-Schwarz inequality, we can write